# SWIFT: A Monotonic, Flux-Form Semi-Lagrangian Tracer Transport Scheme for Flow with Large Courant Numbers

### v1

@@ 5.1.1 Three-Dimensional SWIFT for Density
The first stage is a half step in the vertical. Here the notation is expanded to add a final argument to the flux, $\mathcal{F}(\rho,u,\sigma,\Delta{t})$, which shows the time step used in the flux calculation. If the fourth argument is not shown then it is assumed the time step used is $\Delta{t}$.
@@
$\rho^{z}$ $=\rho^{n}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z}\left(\rho^{n},u^{z},\sigma,\Delta t/2\right),$ (56a)
$\rho_{I}^{x}$ $=\rho^{z}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(\rho^{z}_{A},u^{x}\right),$ (56b)
$\rho_{I}^{y}$ $=\rho^{z}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(\rho^{z}_{A},u^{y}\right),$ (56c)
$\rho^{xy}$ $=\frac{1}{2}\left[\rho_{I}^{x}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(\rho^{x}_{A},u^{y},\sigma^{x}\right)+\rho_{I}^{y}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(\rho_{A}^{y},u^{x},\sigma^{y}\right)\right],$ (56d)
$\rho^{n+1}$ $=\rho^{xy}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z}\left(\rho_{A}^{xy},u^{z},\sigma,\Delta t/2\right),$ (56e)
@@
where $\sigma^{x}$ and $\sigma^{y}$ are given by (33), and
$\sigma^{z}$ $=\sigma-\Delta t\nabla_{z}\cdot\mathcal{F}^{z}(\sigma,u^{z},\sigma,\Delta{t}/2),$ (58a)
$\sigma^{xy}$ $=\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x})-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y}).$ (58b)
@@
$\rho^{n+1}=\rho^{n}-\Delta{t}\nabla_{x}\cdot F^{x}-\Delta{t}\nabla_{y}\cdot F^{y}-\frac{\Delta{t}}{2}\nabla_{z}\cdot F_{1}^{z}+\frac{\Delta{t}}{2}\nabla_{z}\cdot F_{2}^{z},$ (59)
@@
$F_{1}^{z}$ $=\mathcal{F}^{z}(\rho^{n},u^{z},\sigma,\Delta{t}/2),$ (60a)
$F_{2}^{z}$ $=\mathcal{F}^{z}(\rho^{xy}_{A},u^{z},\sigma,\Delta{t}/2),$ (60b)
$F^{x}$ $=\left[\mathcal{F}^{x}(\rho_{A}^{z},u^{x})+\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y})\right]/2,$ (60c)
$F^{y}$ $=\left[\mathcal{F}^{y}(\rho_{A}^{z},u^{y})+\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x})\right]/2.$ (60d)
@@
If $\rho$ is initially a constant, over a whole time step the increment will be proportional to $\bm{\nabla\cdot u}$, thus satisfying the requirement described by Melvin et al. [in review] that density increments should be linear in the divergence.
@@ 5.1.2 Three-Dimensional SWIFT for Tracers
$(\rho m)^{z}$ $=(\rho m)^{n}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z}\left(m^{n},F^{z}_{1},\rho^{z},\Delta t/2\right),$ (62a)
$(\rho m)^{x}$ $=(\rho m)^{z}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(m^{z},F^{x},\rho^{z}\right),$ (62b)
$(\rho m)^{y}$ $=(\rho m)^{z}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(m^{z},F^{y},\rho^{z}\right),$ (62c)
$(\rho m)^{xy}$ $=\frac{1}{2}\left[(\rho m)^{x}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(m^{x},F^{y},\rho^{x}\right)+(\rho m)^{y}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(m^{y},F^{x},\rho^{y}\right)\right],$ (62d)
$(\rho m)^{n+1}$ $=(\rho m)^{xy}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z}\left(m^{xy},F_{2}^{z},\rho^{xy},\Delta t/2\right),$ (62e)
@@ 5.2 The Charney-Phillips Staggering
where $\widetilde{V}_{i}$ is the volume of the shifted mesh cell. Conservative tracer transport therefore conserves the mass that is the global sum of this quantity, whilst solving a transport equation like (2) ensures that this mass is locally-conserved. There are three extra ingredients to obtain the consistency property. Firstly, as in Sections 2.3.2 and 3.2 of this paper, the flux that was used in the transport of $\rho$ must be re-used in the transport of the tracer (but on the shifted mesh). Secondly, the operators to remap fields from the primary mesh to the shifted mesh should satisfy some particular conditions given in Bendall et al. [2023], which guarantee that the processes of transporting a density and converting it to the shifted mesh commute with one another. In this work, we therefore use the remapping operators presented in Bendall et al. [2023], with one amendment: to avoid generating new extrema in the top and bottom half-layers, the transformation between $m$ on the primary and shifted meshes is simply the identity matrix, rather than performing interpolation/extrapolation at the domain boundaries. Finally, the operator to evaluate the flux should satisfy:
$\mathcal{F}(K,F,\rho)=KF,$ (64)
@@ 6 Summary
The benefits of the scheme presented here have seen it become a candidate for the transport scheme used by the Met Office’s new GungHo dynamical core [Melvin et al., in review]. In future work, we will describe how this scheme can be applied to spherical geometries, and in particular cubed-sphere meshes. This scheme can also be combined with the work of Thuburn [2022] to achieve entropy conservation on the Charney-Phillips grid, or with the approach of Brown et al. [2023] to transport tracers which are expressed on a coarser mesh than that used by the dynamical core.
The SWIFT scheme builds upon the work of Lin and Rood [1996], Leonard et al. [1996] and Skamarock [2006] to provide a splitting that retains the properties of the one-dimensional scheme for multi-dimensional transport. The SWIFT scheme, with an appropriate one-dimensional scheme (such as PPM with a limiter), therefore captures the desirable properties of a transport scheme for use in an atmospheric dynamical core, namely: conservation, large time-step-permitting, consistency between tracer transport and density, and monotonicity or positivity.
@@
(https://code.metoffice.gov.uk) upon registration and completion of a software licence. Plotting scripts are stored in a Github repository and will be made available by the authors upon request.
Acknowledgements
The authors would like to acknowledge the frequent discussions with Mohamed Zerroukat, Ben Shipway, Thomas Melvin and Nigel Wood which helped shape this work, and for their feedback on early drafts of the manuscript.
@@ A.1 PPM Coefficients
Let $q_{0}$ and $q_{1}$ be the interpolation of $q$ to cell $i$ facets. Setting $\mathcal{Q}(0)=q_{0}$ and $\mathcal{Q}(1)=q_{1}$ in (4) (i.e. at $\xi=0$ and $\xi=1$ respectively), and setting the integral of the subgrid reconstruction equal to $q_{i}$, gives the PPM coefficients
@@ A.3 PPM Reconstruction
Integrating the PPM coefficients and rearranging gives the PPM facet reconstruction used for the fractional fluxes in the form of (15). As in Section A.1, $q_{0}$ and $q_{1}$ are interpolations of $q$ to the centres of cell facets. Let cell $i$ be the upwind cell from the desired facet, then
@@ B.1 Two-Dimensional Tests
The two-dimensional tests are set up on a doubly periodic planar domain, with $-500\leq x,y\leq 500$ m, so that the domain’s lengths were $L_{x}=1000$ m and $L_{y}=1000$ m. Tests are run to a time of 1000 s.

### v2

@@ 5.1.1 Three-Dimensional SWIFT for Density
The first stage is a half step in the vertical. The corresponding flux calculations use a time step of $\Delta t/2$, and the flux operator is denoted by $\mathcal{F}^{z/2}$.
@@
$\rho^{z}$ $=\rho^{n}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z/2}\left(\rho^{n},u^{z},V\right),$ (61a)
$\rho_{I}^{x}$ $=\rho^{z}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(\rho^{z}_{A},u^{x},\sigma^{z}V\right),$ (61b)
$\rho_{I}^{y}$ $=\rho^{z}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(\rho^{z}_{A},u^{y},\sigma^{z}V\right),$ (61c)
$\rho^{xy}$ $=\frac{1}{2}\left[\rho_{I}^{x}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(\rho^{x}_{A},u^{y},\sigma^{x}V\right)+\rho_{I}^{y}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(\rho_{A}^{y},u^{x},\sigma^{y}V\right)\right],$ (61d)
$\rho^{n+1}$ $=\rho^{xy}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z/2}\left(\rho_{A}^{xy},u^{z},\sigma^{xy}V\right),$ (61e)
@@
where the transported unity fields are given by
$\sigma^{z}$ $=\sigma-\frac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z/2}(\sigma,u^{z},V),$ (63a)
$\sigma^{x}$ $=\sigma^{z}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x},V),$ (63b)
$\sigma^{y}$ $=\sigma^{z}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y},V),$ (63c)
$\sigma^{xy}$ $=\sigma^{z}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x},V)-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y},V).$ (63d)
@@
$\rho^{n+1}=\rho^{n}-\Delta{t}\nabla_{x}\cdot F^{x}-\Delta{t}\nabla_{y}\cdot F^{y}-\frac{\Delta{t}}{2}\nabla_{z}\cdot F_{1}^{z}-\frac{\Delta{t}}{2}\nabla_{z}\cdot F_{2}^{z},$ (64)
@@
$F_{1}^{z}$ $=\mathcal{F}^{z/2}(\rho^{n},u^{z},V),$ (65a)
$F_{2}^{z}$ $=\mathcal{F}^{z/2}(\rho^{xy}_{A},u^{z},V),$ (65b)
$F^{x}$ $=\left[\mathcal{F}^{x}(\rho_{A}^{z},u^{x},V)+\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y}V)\right]/2,$ (65c)
$F^{y}$ $=\left[\mathcal{F}^{y}(\rho_{A}^{z},u^{y},V)+\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x}V)\right]/2.$ (65d)
@@
If $\rho$ is initially a constant, over a whole time step the increment will be proportional to $\bm{\nabla\cdot u}$, thus satisfying the requirement described by Melvin et al. [2024] that density increments should be linear in the divergence.
@@ 5.1.2 Three-Dimensional SWIFT for Tracers
$(\rho m)^{z}$ $=(\rho m)^{n}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z/2}\left(m^{n},F^{z}_{1},\rho^{n}V\right),$ (67a)
$(\rho m)^{x}$ $=(\rho m)^{z}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(m^{z},F^{x},\rho^{z}V\right),$ (67b)
$(\rho m)^{y}$ $=(\rho m)^{z}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(m^{z},F^{y},\rho^{z}V\right),$ (67c)
$(\rho m)^{xy}$ $=\frac{1}{2}\left[(\rho m)^{x}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}\left(m^{x},F^{y},\rho^{x}V\right)+(\rho m)^{y}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}\left(m^{y},F^{x},\rho^{y}V\right)\right],$ (67d)
$(\rho m)^{n+1}$ $=(\rho m)^{xy}-\tfrac{\Delta t}{2}\nabla_{z}\cdot\mathcal{F}^{z/2}\left(m^{xy},F_{2}^{z},\rho^{xy}V\right),$ (67e)
@@ 5.2 The Charney-Phillips Staggering
where $\widetilde{V}_{i}$ is the volume of the shifted mesh cell. Conservative tracer transport therefore conserves the mass that is the global sum of this quantity, whilst solving a transport equation like (2) ensures that this mass is locally-conserved. There are three extra ingredients to obtain the consistency property. Firstly, as in Sections 2.2 and 3.2 of this paper, the flux that was used in the transport of $\rho$ must be re-used in the transport of the tracer (but on the shifted mesh). Secondly, the operators to remap fields from the primary mesh to the shifted mesh should satisfy some particular conditions given in Bendall et al. [2023], which guarantee that the processes of transporting a density and converting it to the shifted mesh commute with one another. In this work, we therefore use the remapping operators presented in Bendall et al. [2023], with one amendment: to avoid generating new extrema in the top and bottom half-layers, the transformation between $m$ on the primary and shifted meshes is simply the identity matrix, rather than performing interpolation/extrapolation at the domain boundaries. Finally, the operator to evaluate the flux should satisfy:
$\mathcal{F}(K,F,\rho V)=KF,$ (69)
@@ 6 Summary
The benefits of the scheme presented here have seen it become a candidate for the transport scheme used by the Met Office’s new GungHo dynamical core [Melvin et al., 2024]. In future work, we will describe how this scheme can be applied to spherical geometries, and in particular cubed-sphere meshes. This scheme can also be combined with the work of Thuburn [2022] to achieve entropy conservation on the Charney-Phillips grid, or with the approach of Brown et al. [2024] to transport tracers which are expressed on a coarser mesh than that used by the dynamical core.
Acknowledgements
The authors would like to acknowledge the frequent discussions with Mohamed Zerroukat, Ben Shipway, Thomas Melvin and Nigel Wood which contributed to shaping this work, and for their feedback on early drafts of the manuscript. The authors are grateful to Hilary Weller whose suggestions helped improve the manuscript, and also to Lucas Harris and the two anonymous referees for their reviews which led to an improved text.
@@ Acknowledgements
(https://code.metoffice.gov.uk) upon registration and completion of a software licence. Plotting scripts are stored in a public Github repository (https://github.com/tommbendall/bendall_kent_swift_2024_plots).
@@ A.1 PPM Coefficients
Let $q_{0}$ and $q_{1}$ be the interpolation of a scalar field $q$ to cell $i$ facets. Setting $\mathcal{Q}(0)=q_{0}$ and $\mathcal{Q}(1)=q_{1}$ in (4) (i.e. at $\xi=0$ and $\xi=1$ respectively), and setting the integral of the subgrid reconstruction equal to $q_{i}$, gives the PPM coefficients
@@ A.3 PPM Reconstruction
Integrating the PPM coefficients and rearranging gives the PPM facet reconstruction used for the fractional fluxes in the form of (15). As in Appendix AA.1, $q_{0}$ and $q_{1}$ are interpolations of $q$ to the centres of cell facets. Let cell $i$ be the upwind cell from the desired facet, then
@@ B.1 Two-Dimensional Tests
The two-dimensional tests are set up on a doubly-periodic planar domain, with $500\leq x,y\leq 500$ m, so that the domain’s lengths were $L_{x}=1000$ m and $L_{y}=1000$ m. Tests are run to a time of 100 s.